the size of trainge abc represent 3 velocity AC magnitude 306 meter per second, corresponde to the velocity of a boat as observed by a stationery tourist on a bridge magnitude 15 meter per second corresponds to the velocity of cylist crossing the bridge, calculate the magnitude of the velocity.
1
Expert's answer
2018-05-18T08:37:08-0400
The magnitude of the velocity is
v=v_1-v_2=306-15=291 m/s.